OpenStudy (anonymous):

is the following always, sometimes or never true 2(8x-4)-2x≤14x+12

OpenStudy (shaik0124):

16x-8-2x<=14x+12 14x-8<=14x+12

OpenStudy (shaik0124):

-8<=12 will be left after subtracting 14x on both sides

OpenStudy (shaik0124):

sometimes true

OpenStudy (shaik0124):

in case of 12>-8 it's true but12=-8 is not true

OpenStudy (shaik0124):

as we know that 12>-8 then its true

OpenStudy (anonymous):

@shaik0124, when would it be *not* true?

OpenStudy (shaik0124):

12 is not eqqual to -8 then its not true

OpenStudy (texaschic101):

2(8x - 4) - 2x <= 14x + 12 16x - 8 - 2x <= 14x + 12 14x - 8 <= 14x + 12 0 <=20 I would say always true sub in numbers and check... lets say x is -2 14(-2) - 8 <= 14(-2) + 12 -28 - 8 <= -28 + 12 - 36 <= - 16 (true) now lets say x = 2 14(2) - 8 <= 14(2) + 12 28 - 8 <= 28 + 12 20 <= 40 (true) now set x = 0 14(0) - 8 <= 14(0) + 12 -8 <= 12 (true)

OpenStudy (anonymous):

@shaik0124, but we're checking to see if the left side is less than OR equal to the right side. It can only be one of these. And since \(-8<12\), it satisfies \(-8\le12\). I'd say always true.

OpenStudy (texaschic101):

I agree

OpenStudy (shaik0124):

done